# Towards Stealthy and Effective Backdoor Attacks on Lane Detection: A Naturalistic Data Poisoning Approach

### v1

@@ Abstract
Deep learning-based lane detection (LD) plays a critical role in autonomous driving and advanced driver assistance systems. However, its vulnerability to backdoor attacks presents a significant security concern. Existing backdoor attack methods on LD often exhibit limited practical utility due to the artificial and conspicuous nature of their triggers. To address this limitation and investigate the impact of more ecologically valid backdoor attacks on lane detection models, we examine the common data poisoning attack and introduce DBALD, a novel diffusion-based data poisoning framework for generating naturalistic backdoor triggers. DBALD comprises two key components: optimal trigger position finding and stealthy trigger generation. Given the insight that attack performance varies depending on the trigger position, we propose a heatmap-based method to identify the optimal trigger location, with gradient analysis to generate attack-specific heatmaps. A region-based editing diffusion process is then applied to synthesize visually plausible triggers within the most susceptible regions identified previously. Furthermore, to ensure scene integrity and a stealthy attack, we introduce two loss strategies: one for preserving lane structure and another for maintaining the consistency of the driving scene. Consequently, compared to existing attack methods, DBALD achieves both a high attack success rate and superior stealthiness. Extensive experiments on 4 mainstream lane detection models show that DBALD exceeds state-of-the-art methods, with an average success rate improvement of +10.87% and significantly enhanced stealthiness. The experimental results highlight significant practical challenges in ensuring model robustness against real-world backdoor threats in lane detection.
@@ Introduction
Deep neural networks (DNNs) have revolutionized autonomous driving systems by enabling robust perception and decision-making capabilities [1, 2, 3, 4]. Among these, lane detection (LD) serves as a cornerstone for safe navigation, as it directly informs path planning and vehicle control. Accurate LD ensures adherence to traffic rules and collision avoidance [5, 6, 7, 8]; however, even minor errors in detection can lead to catastrophic consequences, such as unintended lane departures or collisions. Despite advancements, the reliability of LD models under adversarial conditions remains a critical concern [9, 10, 11]. Recent studies have shown that lane detection models are critically vulnerable to backdoor attacks, especially data poisoning based attack [12, 13, 14]. In such attacks, adversaries need only poison a small portion of the training data (e.g., 1%) by embedding triggers. During inference, these triggers can be activated to manipulate the model’s detection result.
@@
Although several existing works have investigated backdoor threats in the context of LD [15, 16], two fundamental challenges remain unresolved: (i) Suboptimal trigger placement limits attack effectiveness. Current backdoor attack approaches often rely on random or fixed trigger placement strategies, disregarding the critical role of trigger positioning in improving attack success rates. Prior research on backdoor attacks [17, 18] has shown that the spatial location of triggers significantly influences their effectiveness. Indeed, random trigger placement risks positioning triggers in semantically irrelevant or low-attention regions, reducing their memorization by the model during training and thereby diminishing attack efficacy. To illustrate this, Figure 1(a) gives a comparative analysis of fixed triggers injected at three positions (left-top, center, right-bottom), revealing substantial variations in attack success rates across different trigger positions; (ii) Low stealthiness arising from trigger patterns that are detectable by both human observers and forensic detectors Existing backdoor attack methods often exhibit limited stealthiness, primarily due to their reliance on conspicuous trigger designs. Some approaches inject triggers at fixed spatial locations [15], making them visually salient and easily recognizable by human observers. Others superimpose predefined trigger patterns onto input images directly (e.g., Figure 1(c)) [16], often neglecting semantic coherence with the surrounding visual context. Such approaches not only compromise the perceptual realism of the image but also introduce distinguishable artifacts in the frequency domain, making them easily detectable and defendable by forensic detection techniques [19, 20, 21].
@@
Contributions. To address these critical challenges, we present DBALD, a novel and stealthy backdoor attack framework that systematically optimizes both trigger placement and stealth-aware generation for lane detection systems. We first introduce a gradient attention-aware spatial optimization strategy to identify high-sensitivity regions through task-specific heatmap analysis (cf. Figure  1(b)). The heatmap highlights the spatial areas where the LD model exhibits high attention under different tasks (e.g., lane existence prediction or path point regression), guiding the injection of triggers into regions most likely to be memorized during training. Such targeted placement thereby substantially improves the success rate of backdoor activation at inference time. Furthermore, to ensure high stealthiness while maintaining attack performance, building upon recent advances in diffusion models [22, 23, 24], we introduce a semantically-constrained generation pipeline that precisely inpaints contextually appropriate triggers while maintaining strict semantic consistency. Specifically, we introduce two consistency constraints: (i) geometric alignment constraint with existing lane structures, i.e., lane consistency loss, and (ii) photorealistic environmental integration, i.e., environmental consistency loss. These constraints jointly enforce semantic coherence around the edited region, preventing visual artifacts in the editing process. DBALD currently supports three representative attack strategies for the context of LD tasks: lane disappearance attack, lane offset attack, and lane rotation attack. For each attack, we generate corresponding gradient-based heatmaps to guide the optimal trigger placement. Extensive experiments have been conducted over various widely used LD models across both digital and physical simulations. The results show that DBALD achieves a 10.87% improvement in Attack Success Rate (ASR) and a 56.59% improvement in the trigger stealthiness, demonstrating both the effectiveness and the stealthiness.
@@
  • We propose a novel gradient-based attention analysis method tailored to LD tasks that correlates trigger locations with attack effectiveness across different attack strategies;
@@
  • We conduct extensive experiments covering both digital and physical attack scenarios. The experimental results demonstrate that DBALD significantly outperforms existing baselines with an average improvement of +10.87% in attack success rate while maintaining attacking stealthiness.
@@
Lane Detection
As a critical perception task in autonomous driving and advanced driver-assistance systems, lane detection plays a fundamental role in supporting vehicle localization and navigation, obstacle detection and avoidance, automatic parking, etc [25, 26, 27]. Early approaches primarily rely on handcrafted features [28, 29] and geometric priors [30, 31], but are sensitive to illumination variations and complex road conditions. The last decade has witnessed the rapid development of deep learning-based LD methods, which have significantly improved the accuracy and robustness under diverse driving scenarios. They can be separated to anchor-based [32, 33], segmentation-based [34, 35], keypoint-based [36, 37] and parametric curved-based [38, 39] methods, with the first two being the most mainstream approaches [25]. Anchor-based methods detect lanes by classifying and regressing predefined anchors placed across the image. Their efficiency and strong real-time performance have made them a popular choice in modern LD systems. Segmentation-based methods regard LD as an image segmentation problem, leveraging deep neural networks to extract high-level semantic features for distinguishing lane regions from the background. By generating dense, per-pixel predictions, these methods excel at handling complex road structures and challenging visual conditions. Despite their excellent performance, these LD models are exposed to significant security vulnerabilities, and in this work, we demonstrate how backdoor attacks can maliciously compromise their reliability.
Backdoor Attacks
@@
In backdoor attacks, the model is trained using a combination of clean samples and a small proportion of poisoned samples embedded with a specific trigger [13, 40, 41]. While the model behaves normally on clean inputs, it predicts a wrong output whenever the trigger is present in the input. Such attacks are difficult to detect, since the performance on standard validation data remains unaffected, and only the attacker who implants the trigger knows how to activate the backdoor. In lane detection scenarios, once the detector misbehaves, it can lead to critical driving errors, such as unintended lane departures or incorrect path planning, posing serious risks to the safety and reliability of autonomous vehicles [15, 16]. Despite increasing research efforts on backdoor attacks in image classification [42, 43, 44, 45] and LiDAR perception [46, 47, 48], backdoor vulnerabilities in lane detection remain largely unexplored. Han et al. [15] were the pioneering researchers to investigate this field. They proposed poison-annotation and clean-annotation methods to generate poisoned samples that could effectively compromise lane detection models. However, their work was done on static scenes with unchanging environmental contexts, which severely restricts their applicability in the dynamic environments of autonomous driving. To solve this problem, Zhang et al. [16] proposed Badlane which incorporates an amorphous trigger pattern inspired by mud spots and a meta-learning framework to generate triggers that adapt to varying environmental conditions, thereby enhancing the robustness of the attack in dynamic driving scenarios. However, Badlane incurs higher computational overhead and its trigger pattern remains visually conspicuous, increasing the risk of being detected. Moreover, BadLANE’s random trigger placement often occurs in low-attention regions, reducing the model’s ability to learn backdoor behavior and thereby lowering ASR. Recent efforts such as PEFTGuard [49] highlight the vulnerability of PEFT adapters to backdoor injection and propose effective detection techniques. Similarly, stealthy attacks on multimodal models [50] demonstrate that semantically aligned yet inconspicuous triggers can effectively bypass detection. These insights motivate more naturalistic and stealth-aware attacks for lane detection. In this paper, we propose a novel diffusion-based backdoor injection framework that improves both attack stealthiness and effectiveness. By leveraging the attention maps, we identify the key regions where the LD model focuses most during training to guide the trigger placement. A diffusion-based trigger generation method further allows DBALD to improve the attack stealthiness while maintaining the ASR.
@@
[…]
We inherit the definition of the backdoor trigger pattern given in [16]. Specifically, a trigger $\mathcal{T}$ is a set of pixels denoted as follows:
@@
[…]
@@
We assume that the attacker has access to the original training dataset of the lane detection model and also has prior knowledge of the LD model that the victim will deploy. The attack seeks to poison the training dataset so that the resulting model performs normally on benign inputs but generates erroneous predictions when a specific trigger pattern is present.
@@
In this work, we consider three mainstream attack strategies of lane detection models: Lane Offset Attack (LOA), Lane Disappearance Attack (LDA), and Lane Rotation Attack (LRA) [16, 15].
Lane Offset Attack: LOA simulates a scenario where all lane positions are offset by a fixed number of pixels $\beta$ from their true positions. This misalignment causes the vehicle to deviate from the correct lane, potentially resulting in unsafe driving conditions. Given an image $x$ and its ground truth label $\tilde{g}=[(\tilde{l}_{i},\tilde{e}_{i})\mid 1\leq i\leq n]$ with $\tilde{l}_{i}=\{\tilde{p}^{i}_{1},\ldots,\tilde{p}^{i}_{m}\}$, the attacked prediction result by LOA can be denoted as $g=[(l_{i},e_{i})\mid i\in[n]]$ such that: $\forall i\in[n]$,
@@
Backdoor Trigger. We assume that the attacker aims to use practical or real-world physical triggers to mislead the lane detection models. These triggers could take the form of, for example, a cone with a unique shape or mud with a characteristic brown color.
DBALD Design
@@
The working pipeline of DBALD is illustrated in Figure 3, consisting of three major stages: trigger position candidates generation, trigger optimal position finding, stealthy trigger generation.
Trigger Position Candidates Generation. As part of the data preprocessing step, road-region segmentation is performed on the original training samples using an existing segmentation method [51] to identify realistic candidate regions for trigger placement. Specifically, for the lane detection task, we select the road region as the primary candidate area, and road-region constraints are then derived to form a pool of potential trigger positions. To refine this pool into concrete position candidates, we employ a sliding window module over the SAM-segmented road region, using a fixed stride to generate a discrete set of valid trigger position candidates.
@@
Trigger Generation. Once the optimal trigger position is determined, a masked image is created by masking the trigger region, and the surrounding environmental context of the trigger position is extracted as the ground truth for the trigger’s environment. This masked image is then fed into a carefully designed mask-guided diffusion model, which inpaints the trigger in a stealthy manner. To ensure seamless integration, the diffusion model is refined in the latent space through the incorporation of realness-preserving loss terms. These loss terms leverage the ground truth of the trigger’s surrounding environment context to guide the generation process, ensuring that the synthesized trigger appears visually plausible and blends naturally into the scene. Finally, the poisoned samples, containing the stealthily embedded triggers, are injected back into the training dataset. This would compromise the integrity of the lane detection model trained on the dataset, enabling the triggers to be effectively activated during the inference stage.
@@
Prior studies [17, 18] have demonstrated that embedding triggers in regions with higher attention enhances the ability of DNN models to memorize trigger patterns during training, thereby improving the attack’s effectiveness during inference. However, conventional attention mechanisms, such as Grad-CAM [52], are predominantly tailored for pure classification tasks. In contrast, lane detection models inherently involve hybrid tasks, including lane existence and lane coordinate predictions, which require a more specialized approach. To address this gap and extract high-attention information relevant to our task, we propose a gradient-based attention map to identify optimal trigger positions.
@@
$M_{\text{grad}}(i,j)=\sum_{k=1}^{CH}\left|\frac{\partial L}{\partial x_{i,j,k}}\right|,\quad\text{for input image }x\in\mathbb{R}^{H\times W\times CH},$
[…]
Moreover, to align with different attack strategies, we design strategy-specific heatmaps by leveraging gradient information from task-relevant loss functions. Since different LD architectures are trained with different types of loss functions, we generate gradient attention heatmaps accordingly. Specifically, anchor-based methods (e.g., LaneATT [32], AdNet [33]) typically use focal loss for lane existence prediction (used in LDA) and L1/GLIoU loss for coordinate regression (used in LOA and LRA). Segmentation-based methods (e.g., SCNN [35], RESA [34]) rely on cross-entropy loss for classification and binary cross-entropy loss for regression, respectively. More details of these two types of LD can be viewed on [53].
Figure 4: From left to right: the original image, the image with the mask, the ground truth of surrounding environmental objects, the attacked image generated from vanilla mask-guided diffusion, and the generation result from our method. The bottom row demonstrates that, in the absence of the proposed lane and environment loss terms, the diffusion model may produce inconsistent or unrealistic structures, such as deformed lane markings (above) and incomplete vehicles (below). In contrast, our method effectively preserves contextual integrity, ensuring more accurate and visually coherent reconstructions.
Building on these foundations, we compute the optimal trigger position as follows. Given an input image and a specific attack strategy, we first generate a strategy-specific heatmap by leveraging the gradients of the corresponding loss functions. For example, in the case of LDA, classification loss gradients are used to suppress lane existence confidence, while for LOA or LRA, regression loss gradients are utilized to distort geometric features. Next, we identify high-sensitivity regions within the road areas segmented by SAM2 [51]. To pinpoint the most suitable locations for trigger placement, we employ a sliding window algorithm that identifies rectangular patches with maximal gradient responses within these high-sensitivity regions. This approach typically generates approximately 100400100$\sim$ 400100 ∼ 400 potential trigger candidates. By integrating the previously computed heatmap gradient results, we evaluate these candidates to determine the optimal trigger position with the highest attention value. The entire procedure ensures that the triggers are embedded in positions that exert the greatest influence on the targeted tasks, thereby maximizing the attack’s efficacy.
@@
Based on the regions identified as optimal for trigger placement in Section 4.2, we adopt a region-based diffusion editing framework inspired by UltraEdit [22], which facilitates precise and controlled injection of backdoor triggers within specific mask regions. Specifically, this mask-guided diffusion editing selectively modifies only the masked region, ensuring the generation of visually plausible and realistic backdoor triggers while leaving other areas of the image as unaffected as possible.
@@
[…]
@@
However, the inherent attention mechanisms of diffusion models may still inadvertently distort crucial visual elements, such as lane markings or surrounding traffic objects located adjacent to the mask boundaries. To mitigate this issue and improve the stealthiness of the attack, we introduce two additional loss terms, $L_{\text{lane}}$ and $L_{\text{env}}$, into the diffusion process: $L_{\text{lane}}&=\text{MSE}\left(\text{gen\_img}\odot\text{lane\_mask},\ \text{clean\_img}\odot\text{lane\_mask}\right),\\ L_{\text{env}}&=\text{SSIM}\left(\text{gen\_img}\odot\text{env\_mask},\ \text{clean\_img}\odot\text{env\_mask}\right)\end{aligned}$
$L_{\text{lane}}$ is designed to guide the generated image to closely align with the original (clean) image within the lane regions, as specified by the binary lane_mask. By minimizing this mean squared error, we ensure that the lane structures remain visually intact and consistent, thereby reducing potential artifacts that could otherwise compromise the realism of the scene. In contrast, $L_{\text{env}}$ employs structural similarity (SSIM) to preserve the visual integrity and realism of surrounding traffic elements and environmental structures. This ensures that the modifications made during the diffusion process do not distort the contextual coherence of the scene, maintaining a high level of stealthiness in the attack.
We note that the incorporation of these two loss terms is crucial for enhancing the stealthiness of the generated triggers by effectively mitigating potential artifacts introduced by the attention mechanisms of diffusion models. It indeed helps us generate poisoned samples that are seamless and natural-looking. As illustrated in Figure 4, the absence of these loss terms (i.e., using vanilla mask-guided diffusion) typically leads to noticeable discrepancies in the reconstructed content, particularly in regions adjacent to the mask. In contrast, the proposed method successfully preserves critical environmental cues, producing reconstructions that are visually coherent and semantically faithful to the original scene.
@@
ACC ASR(Diff. strategies) ASR ACC ASR(Diff. strategies) ASR
Vanilla: Benign (Poisoned) Infected LDA  LRA  LOA Avg. Vanilla: Benign (Poisoned) Infected   LDA   LOA Avg.
LaneATT BadNets 75.01 (74.97) 72.79 49.27   43.37   47.36 46.67 96.81 (95.92) 95.53   45.62   75.63 60.63
@@
DBALD 74.20 81.65   73.45   74.19 77.23 95.72   90.04   91.63 90.84
ADNet BadNets 85.51 (84.94) 79.60 50.13   54.33   61.39 55.28 96.87 (96.49) 95.46   50.34   52.98 51.66

### v2

@@ Abstract
Deep learning-based lane detection (LD) plays a critical role in autonomous driving and advanced driver assistance systems. However, its vulnerability to backdoor attacks presents a significant security concern. Existing backdoor attack methods on LD often exhibit limited practical utility due to the artificial and conspicuous nature of their triggers. To address this limitation and investigate the impact of more ecologically valid backdoor attacks on lane detection models, we examine the common data poisoning attack and introduce DBALD, a novel diffusion-based data poisoning framework for generating naturalistic backdoor triggers. DBALD comprises two key components: optimal trigger position finding and stealthy trigger generation. Given the insight that attack performance varies depending on the trigger position, we propose a heatmap-based method to identify the optimal trigger location, with gradient analysis to generate attack-specific heatmaps. A region-based editing diffusion process is then applied to synthesize visually plausible triggers within the most susceptible regions identified previously. Furthermore, to ensure scene integrity and a stealthy attack, we introduce two loss strategies: one for preserving lane structure and another for maintaining the consistency of the driving scene. Consequently, compared to existing attack methods, DBALD achieves both a high attack success rate and superior stealthiness. Extensive experiments on 4 mainstream lane detection models show that DBALD exceeds state-of-the-art methods, with an average success rate improvement of +10.87% and significantly enhanced stealthiness. The experimental results highlight significant practical challenges in ensuring model robustness against real-world backdoor threats in lane detection. Our data and demos are available at https://sites.google.com/view/dbald.
@@ Introduction
Deep neural networks (DNNs) have revolutionized autonomous driving systems by enabling robust perception and decision-making capabilities [bachute2021autonomous, cui2021deep, mozaffari2020deep, kuutti2020survey]. Among these, lane detection (LD) serves as a cornerstone for safe navigation, as it directly informs path planning and vehicle control. Accurate LD ensures adherence to traffic rules and collision avoidance [gehrig2007collision, xu2025vlm, hu2023planning, amaradi2016lane]; however, even minor errors in detection can lead to catastrophic consequences, such as unintended lane departures or collisions. Despite advancements, the reliability of LD models under adversarial conditions remains a critical concern [muhammad2020deep, xiaomi_su7_crash, grigorescu2020survey]. Recent studies have shown that lane detection models are critically vulnerable to backdoor attacks, especially data poisoning based attacks [wu2022backdoorbench, li2022backdoor, fang2022backdoor]. In such attacks, adversaries need only poison a small portion of the training data (e.g., 10%) by embedding triggers. During inference, these triggers can then be activated to manipulate the model’s detection result.
@@
Although several existing works have investigated backdoor threats in the context of LD [han2022physical, zhang2024towards], two fundamental challenges remain unresolved: (i) Suboptimal trigger placement limits attack effectiveness. Current backdoor attack approaches often rely on random or fixed trigger placement strategies, disregarding the critical role of trigger positioning in improving attack success rates. Prior research on backdoor attacks [wenger2021backdoor, abad2023sok] has shown that the spatial location of triggers significantly influences their effectiveness. Indeed, random trigger placement risks positioning triggers in semantically irrelevant or low-attention regions, reducing their memorization by the model during training and thereby diminishing attack efficacy (e.g., Figure 1(a)). (ii) Low stealthiness arising from trigger patterns that are detectable by both human observers and forensic detectors Existing backdoor attack methods often exhibit limited stealthiness, primarily due to their reliance on conspicuous trigger designs. Some approaches inject triggers at fixed spatial locations [han2022physical], making them visually salient and easily recognizable by human observers. Others superimpose predefined trigger patterns onto input images directly (e.g., Figure 1(c)) [zhang2024towards], neglecting semantic coherence with the surrounding visual context. Such approaches not only compromise the perceptual realism of the image but also introduce distinguishable artifacts in the frequency domain, making them easily detectable by forensic detection techniques [ojha2023towards, tan2023learning, stamm2010forensic].
@@
Contributions. To address the above critical challenges, we propose DBALD, a novel and stealthy backdoor attack framework for lane detection systems that optimizes both trigger placement and generation. First, we introduce a gradient attention-based spatial strategy that leverages task-specific heatmaps (e.g., for lane existence or lane point regression) to identify high-sensitivity regions for trigger injection, significantly boosting attack success rates.(e.g., Figure 1(b,d)) To enhance stealthiness, we employ a diffusion-based inpainting pipeline guided by two semantic constraints: (i) lane consistency loss for geometric alignment with existing lane structures, and (ii) environmental consistency loss for photorealistic scene integration. These constraints ensure visually coherent and contextually plausible triggers. DBALD supports three backdoor attack strategies: lane disappearance, offset, and rotation attacks. Each attack strategy is guided by its own optimized heatmap. Extensive evaluations across multiple LD models and settings show that DBALD improves Attack Success Rate (ASR) by 10.87% and stealthiness by 56.59% over prior methods.
@@
  • We propose a novel gradient-based attention analysis method tailored to LD tasks that correlates trigger locations with ASR across different attack strategies;
@@
  • We conduct extensive experiments covering both digital and physical attack scenarios. The results demonstrate that DBALD significantly outperforms existing baselines with an average improvement of +10.87% in attack success rate while maintaining attacking stealthiness.
@@
Lane Detection. As a key perception task in autonomous driving, lane detection supports vehicle navigation and obstacle avoidance [bi2025lane, ma2024monocular]. Traditional methods based on geometric priors [gao2010sobel, dorj2020kalman] suffer precision limits under complex scenes. Recent deep learning-based approaches greatly improve robustness and can be categorized into anchor-based [tabelini2021keep, xiao2023adnet], segmentation-based [zheng2021resa, pan2018spatial], keypoint-based [wang2022keypoint], and curve-based [feng2022rethinking]. Among them, anchor-based and segmentation-based methods dominate due to their accuracy and efficiency. However, despite their success, these models remain vulnerable to adversarial threats. In this work, we show how backdoor attacks can compromise the reliability of modern lane detection systems.
@@
Backdoor Attacks. In backdoor attacks, the model is trained using a combination of clean samples and a small proportion of poisoned samples embedded with a specific trigger [li2022backdoor, liu2018trojaning, chen2017targeted, guo20256dattack, zhang2025backdoor, zhong2025backdoor]. While the model behaves normally on clean inputs, it predicts a wrong output whenever the trigger is present in the input. Such attacks are difficult to detect, since the performance on clean validation data remains unaffected, and only the attacker who implants the trigger knows how to activate the backdoor. In lane detection scenarios, once the detector predicts incorrectly, it would lead to critical driving errors, posing serious risks to the safety and reliability of autonomous vehicles [han2022physical, zhang2024towards]. Despite increasing research efforts on backdoor attacks in image classification [pourkeshavarz2024adversarial, saha2020hidden, doan2021lira, li2021invisible, ni2024physical] and LiDAR perception [zhang2022towards, li2023towards, xiang2021backdoor], backdoor vulnerabilities in lane detection remain largely unexplored. Han et al. [han2022physical] introduced annotation-based poisoning but lacked adaptability to dynamic scenes. Zhang et al.’s Badlane [zhang2024towards] improved robustness with amorphous triggers and meta-learning, yet suffered from visual conspicuity and suboptimal trigger placement. In this paper, we propose DBALD, a diffusion-based backdoor framework that improves stealth and effectiveness by placing triggers in high-attention regions and generating them via context-aware diffusion inpainting.
@@
Given an LD model $f_{\theta}$ parameterized by $\theta$ and an input image $x$, the model predicts an output $g$ such that $f_{\theta}(x)\to g$. Specifically, $g$ consists of a series of lane predictions represented as $g=[(l_{1},e_{1}),(l_{2},e_{2}),\dots,(l_{n},e_{n})]$, where $l_{i}=\{p^{i}_{1},p^{i}_{2},\ldots,p^{i}_{m}\}$ denotes a sequence of point coordinates of the $i$-th lane, $e_{i}\in\{0,1\}$ is a binary indicator representing whether the corresponding lane exists, $n$ denotes the lane number and $m$ denotes the coordinate number. The ground truth label of $x$ is denoted as $\tilde{g}=[(\tilde{l}_{1},\tilde{e}_{1}),(\tilde{l}_{2},\tilde{e}_{2}),\dots,(\tilde{l}_{n},\tilde{e}_{n})]$. Typically, for a well-trained LD model, it is expected to predict the lane coordinates and the existence indicators that are equal to the ground truth $\tilde{g}$, i.e., $f_{\theta}(x)=\tilde{g}$. We use $[n]$ to denote $\{1,2,\ldots,n\}$.
We inherit the definition of the backdoor trigger pattern given in [zhang2024towards]. Specifically, a trigger $\mathcal{T}$ is a set of pixels denoted as follows:
@@
where $q[(w,h),c]$ denotes a pixel located at $(w,h)$ with color $c$, $R$ is the trigger region, and $C$ is the color space.
@@
We assume that the attacker has access to the original training dataset of the LD model and also has prior knowledge of the LD model that the victim will deploy. The attack seeks to poison the training dataset so that the resulting model performs normally on benign inputs but generates erroneous predictions when a specific trigger pattern is present.
@@
In this work, we consider three mainstream attack strategies of LD models: Lane Offset Attack (LOA), Lane Disappearance Attack (LDA), and Lane Rotation Attack (LRA) [zhang2024towards, han2022physical].
Lane Offset Attack: LOA simulates a scenario where all lane positions are offset by a fixed number of pixels $\beta$ from their true positions. To be specific, Given an image $x$ and its ground truth label $\tilde{g}=[(\tilde{l}_{i},\tilde{e}_{i})\mid 1\leq i\leq n]$ with $\tilde{l}_{i}=\{\tilde{p}^{i}_{1},\ldots,\tilde{p}^{i}_{m}\}$, the attacked prediction result by LOA can be denoted as $g=[(l_{i},e_{i})\mid i\in[n]]$ such that: $\forall i\in[n]$,
@@
Backdoor Trigger. We assume that the attacker aims to use practical physical triggers to mislead the LD models. These triggers could take the form of, for example, a cone with distinctive shapes or colors, or mud with a characteristic color following existing work [zhang2024towards, han2022physical].
Methodology
@@
As shown in Figure 3, DBALD consists of three stages: trigger position candidates generation, optimal position selection, and stealthy trigger generation.
Trigger Position Candidates Generation. We segment the road region using SAM2 [ravi2024sam2] to identify plausible areas for trigger placement. A sliding window is then applied over the segmented mask to generate a discrete set of position candidates.
@@
Trigger Generation. Once the optimal trigger position is determined, a masked image is created by masking the trigger region, and the surrounding environmental context of the trigger position is extracted as the ground truth for the trigger’s environment. This masked image is then fed into a carefully designed mask-guided diffusion model, which inpaints the trigger in a stealthy manner. To ensure seamless integration, the diffusion model is refined in the latent space through the incorporation of realness-preserving loss terms. These loss terms leverage the ground truth of the trigger’s surrounding environment context to guide the generation process, ensuring that the synthesized trigger appears visually plausible and blends naturally into the scene. Finally, the poisoned samples, containing the stealthily embedded triggers, are injected back into the training dataset. This would compromise the integrity of the LD model trained on the dataset, enabling the triggers to be effectively activated during the inference stage.
@@
Prior studies [wenger2021backdoor, abad2023sok] have demonstrated that embedding triggers in regions with higher attention enhances the ability of DNN models to memorize trigger patterns during training, thereby improving the attack’s effectiveness during inference. However, conventional attention mechanisms, such as Grad-CAM [selvaraju2017grad], are predominantly tailored for pure classification tasks. In contrast, LD models inherently involve hybrid tasks, including lane existence and lane coordinate predictions, which require a more specialized approach. To address this gap and extract high-attention information relevant to our task, we propose a gradient-based attention map to identify optimal trigger positions.
@@
$M_{\text{grad}}(i,j)=\sum_{k=1}^{CH}\left|\frac{\partial L}{\partial x_{i,j,k}}\right|,\text{for input image }x\in\mathbb{R}^{H\times W\times CH}$
where $L$ denotes the task-specific loss function, $(i,j)$ denotes the spatial position in the image, $k$ is the index of the channel dimension, and $CH$ represents the total number of channels. The resulting attention value $M_{\text{grad}}(i,j)$ is computed by summing the absolute gradient magnitudes across all channels at the given location, thereby capturing the sensitivity of that location with respect to the loss function.
Moreover, to align with different attack strategies, we design strategy-specific heatmaps by leveraging gradient information from task-relevant loss functions. Since different LD models are trained with different types of loss functions, we generate gradient attention heatmaps accordingly. Specifically, anchor-based methods (e.g., LaneATT [tabelini2021keep], AdNet [xiao2023adnet]) typically use focal loss for lane existence prediction (used in LDA) and L1/GLIoU loss for coordinate regression (used in LOA and LRA). Segmentation-based methods (e.g., SCNN [pan2018spatial], RESA [zheng2021resa]) rely on cross-entropy loss for classification and binary cross-entropy loss for regression, respectively. More details of these two types of LD are in supplementary materials.
Figure 4: The bottom row demonstrates that, in the absence of the proposed lane and environment loss terms, the diffusion model may produce inconsistent or unrealistic structures, such as deformed lane markings (above) and incomplete vehicles (below). In contrast, our method effectively preserves contextual integrity, ensuring more accurate and visually coherent reconstructions.
Building on these foundations, we compute the optimal trigger position as follows. Given an input image and a specific attack strategy, we first generate a strategy-specific heatmap by leveraging the gradients of the corresponding loss functions. For example, in the case of LDA, classification loss gradients are used to suppress lane existence confidence, while for LOA or LRA, regression loss gradients are utilized to distort geometric features. Next, we identify high-sensitivity regions within the road areas segmented by SAM2 [ravi2024sam2]. To pinpoint the most suitable locations for trigger placement, we employ a sliding window algorithm that identifies rectangular patches with maximal gradient responses within these high-sensitivity regions. This approach typically generates approximately $100\sim 400$ potential trigger candidates. By integrating the previously computed heatmap gradient results, we evaluate these candidates to determine the optimal trigger position with the highest attention value. The procedure ensures triggers are embedded in positions that exert the greatest influence on the targeted tasks, thereby maximizing the attack’s efficacy.
@@
Based on the regions identified as optimal for trigger placement in Section 4.2, we adopt a region-based diffusion editing framework inspired by UltraEdit [zhao2024ultraedit], which facilitates precise and controlled injection of backdoor triggers within specific mask regions. Specifically, this mask-guided diffusion editing selectively modifies only the masked region, ensuring the generation of visually plausible and realistic backdoor triggers while leaving other areas of the image as unaffected as possible.
@@
where $\odot$ denotes element-wise multiplication, and $M\in\{0,1\}^{H\times W\times C}$ is the binary mask that specifies the trigger region, $M(i,j,c)=1$ demonstrates that the pixel at position $(i,j)$ in channel $c$ is designated for modification. $D_{M}(z_{t})$ refers to the denoising function applied to $z_{t}$, with updates restricted to the masked region defined by $M$. When the diffusion step is even, the diffusion process is confined strictly to the specified masked region, leaving the latent encoding of the original image unchanged outside the mask. Conversely, when the diffusion step is odd, standard diffusion is applied across the entire image. Due to the preservation of the original latent encoding during the even steps, regions outside the mask remain largely unaffected.
@@
However, the inherent attention mechanisms of diffusion models still inadvertently distort crucial visual elements, such as lane markings or surrounding traffic objects located adjacent to the mask boundaries. To mitigate this issue and improve the stealthiness of the attack, we introduce two additional loss terms, $L_{\text{lane}}$ and $L_{\text{env}}$, into the diffusion process:
$\begin{aligned} L_{\text{lane}}&=\text{MSE}\left(\text{gen\_img}\odot\text{lane\_mask},\ \text{clean\_img}\odot\text{lane\_mask}\right)\\ L_{\text{env}}&=\text{SSIM}\left(\text{gen\_img}\odot\text{env\_mask},\ \text{clean\_img}\odot\text{env\_mask}\right)\end{aligned}$
$L_{\text{lane}}$ is designed to guide the generated image to closely align with the original (clean) image within the lane regions, as specified by the binary lane_mask. By minimizing this mean squared error, we ensure that the lane structures remain visually intact and consistent, thereby reducing potential artifacts that could otherwise compromise the realism of the scene. $L_{\text{env}}$, on the other hand, employs structural similarity (SSIM) to preserve the visual integrity and realism of surrounding traffic elements and environmental structures. This ensures that the modifications made during the diffusion process do not distort the contextual coherence of the scene, maintaining a high level of stealthiness in the attack.
Incorporating these two loss terms is essential for improving the stealthiness of the generated triggers, as they effectively suppress the artifacts typically introduced by diffusion models. This design choice enables the creation of poisoned samples that appear seamless and natural. As illustrated in Figure 4, removing these loss terms (i.e., adopting vanilla mask-guided diffusion) often leads to noticeable inconsistencies in the reconstructed content, especially in areas adjacent to the mask. In contrast, our proposed method preserves critical environmental cues and yields reconstructions that remain visually coherent and semantically faithful to the original scene.
Clarification on Trigger Diversity.
To quantify trigger diversity, we compute pairwise perceptual similarity using LPIPS (Learned Perceptual Image Patch Similarity) scores [zhang2018unreasonable] and adopt a threshold of 0.15, following prior work [hou2022perceptual]. This threshold guides the construction of practical triggers by ensuring that each trigger retains a distinctive visual signature while remaining sufficiently dissimilar from benign patterns that may, albeit rarely, appear in the dataset. Consequently, it preserves the clean-test accuracy of the poisoned model while maintaining a high ASR.
@@
ACC on clean images ASR (Diff. strategies) ASR ACC on clean images ASR (Diff. strategies) ASR
Benign Poisoned LDA  LRA  LOA Avg. Benign Poisoned   LDA   LOA Avg.
LaneATT BadNets 72.79 49.27   43.37   47.36 46.67 95.53   45.62   75.63 60.63
@@
DBALD 74.20 81.65   73.45   74.19 76.41 95.72   90.04   91.63 90.84
ADNet BadNets 85.51 79.60 50.13   54.33   61.39 55.28 96.87 95.46   50.34   52.98 51.66